An RC circuit have a 9 V source and lies at a 54-seconds of time constant.

Then, find the voltage across the Capacitor after 153.6 seconds of charging.


1
Expert's answer
2022-04-03T16:51:01-0400
"V_C(t)=V_s\left(1-e^{-t/\tau}\right)"

"V_C(t)=9\:{\rm V}\left(1-e^{-153.6/54}\right)=8.48\:\rm V"